# Towards Open Federated Learning Platforms: Survey and Vision from Technical and Legal Perspectives

### 18 Jan 2024 v2

@@ 5. Contract-based Federated Learning
DeDES (Wang et al., 2023b) Model Subset Model
@@
FedProto (Tan et al., 2022a; Michieli and Ozay, 2021) Class
@@ 5.1. How to Design ML Microtasks
When we come to the scenario of contract-based FL, we first need to define how to design ML microtasks for crowd workers. Unfortunately, previous collaborative ML studies (Li et al., 2021c; Nguyen et al., 2021) rarely discuss this question because they default to assuming that all employers or workers have the same resource type to fit into their modeling frameworks. On the contrary, in contract-based FL, we leave the freedom of design microtasks and choosing modeling method to the employers.
@@
The answer to this question depends on what resources the employers have and what tasks the workers can perform. For example, in horizontal FL, the server should provide an initial model, and clients should have training data under the same feature space and provide computational power. However, in vertical FL, the clients should have training data under the same sample space (i.e. ID space). This difference determines which modeling framework employers can adopt. Following this idea, we provide comparisons of typical FL, distributed ML, and blockchain-based modeling frameworks in Table 9. For simplicity, we introduce the Platform between Employer and Worker in some frameworks to make it adaptable for contract-based FL. You can quickly restore the original structure by merging Employer and Platform columns. To represent the transmission of resources, we mark resources provided by each group of entities with different colors. The computational resource used for training is marked with $\bigstar$. By this way, we can conveniently find the feasible and privacy-preserving modeling frameworks for collaborative learning. Here, we briefly introduce these frameworks group by scenarios [^22] From the viewpoint of Employers..
5.1.1. Train from scratch with workers’ private data
@@
There are four frameworks that support this scenario. The first is Centralized ML, which requires workers to upload their data to a cloud platform for modeling. However, this approach is not feasible for FL scenarios due to privacy concerns. The second is FedAvg (McMahan et al., 2017), which requires workers to upload computed gradients upon the scratch model and their private data (horizontally split). Computing resources need to be provided by workers as well. The third is SplitNN (Vepakomma et al., 2019), where each worker trains the cut layers of the model using their vertically split data. The hidden representations from these layers are uploaded to the platform for the training of the remaining layers. The last and distinctive is a Web3 system named Ocean (McConaghy, 2022). It enables workers to upload only the metadata of their local dataset to a blockchain-based platform to attract buyers. Interested buyers then need to purchase datatokens to access the local dataset and deploy their modeling algorithm. To ensure the privacy and IP security of workers, only trusted algorithms are allowed, and only model predictions will be sent to buyers.
5.1.2. Ensemble workers’ private model
As we presented in Section 4.3.4, ensembling is an effective method to integrate knowledge, and the distributed version is proposed by DMoE (Ryabinin and Gusev, 2020). It employs a Distributed Hash Table (DHT) to store metadata and worker node statuses, constructing a decentralized expert network. To make inferences using this network, each model user must train a local gating network to select a subset of experts tailored to their input. No additional training required, Wang et al. purposed DeDES (Wang et al., 2023b), which select a diverse subsect of weak models from population and make inference by voting. The unique advantage of ensembling lies in its inherent support for model heterogeneity, and its integration is extensible. This capability enables the establishment of a cooperative network while maintaining flexibility and availability.
5.1.3. Train from fundation model
@@
Foundation models, trained on larger-scale data with robust generalization abilities across various downstream tasks, serve as a solid foundation for collaborative training. Yuan et al. proposed DT-FM (Yuan et al., 2022) to establish an effective geo-distributed learning system across 8 regions for training the language model GPT-3. Similarly, Borzunov et al.(Borzunov et al., 2022) recruited volunteer nodes to train a transformer model over the Internet. To enhance system robustness, Moshpit SGD (Ryabinin et al., 2021) divides worker nodes into small, independent groups to ensure that all-reduce is not affected by the failure of a single participant. Additionally, VC-ASGD(Atre et al., 2021) divides the deep learning training job into asynchronous parameter update subtasks to improve scalability. In these cases, employers are required to supply the initial model and training data needed to launch training subtasks. This training paradigm is particularly suitable for worker nodes that possess computational and communication resources but lack their own training data.
@@
Instead of initiating training from scratch and consuming substantial computational resources, we can also collaboratively adapt these foundation models to specific local tasks based on relatively restricted hardware and data size through fine-tuning. Recently, Woisetschläger et al. (Woisetschläger et al., 2023) have explored the opportunity to fine-tune large language models, such as FLAN-T5, by edge devices. Simultaneously, FS-LLM (Kuang et al., 2023) employs Parameter-Efficient Fine-Tuning (PEFT) methods for federated fine-tuning of LLaMA-7B, minimizing communication and computation costs. FATE-LLM (Fan et al., 2023) offers a solution called FedHeteroLLM, leveraging KD to train a mentee model from its local pre-trained LLM for federated aggregation. To alleviate the computational burden associated with backpropagation-based optimization methods, FedKSeed (Qin et al., 2023) employs zeroth-order optimization (ZOO). Only seed and scalar gradients need to be transmitted for federated fine-tuning.
5.1.4. Representations Sharing
@@
Another popular scheme for organizing ML microtasks involves sharing the learned representations of workers. For instance, the use of class-conditional average of last hidden layer activations (aka. Prototypes) can enhance class discrimination across different clients (Aketi and Roy, 2024; Tan et al., 2022a; Michieli and Ozay, 2021; Berdoz et al., 2022). To configure the microtasks, employers only need to set the target class, after which the platform can collect and distribute per-class prototypes among workers. With no central server required, CCL (Aketi and Roy, 2024) presents a decentralized learning approach based on cross-workers prototypes sharing.
5.2. Decentralization and Monetization
Parallel to contract-based FL, several efforts are underway to build decentralized AI platforms through Web3-based techniques (Guo et al., 2023). For instance, Blythman et al. proposed decentralized AI Hubs (Blythman et al., 2022), empowered by DAOs (e.g., IPFS (Benet, 2014), Ocean (McConaghy, 2022), OpenMined), to run ML activities in a trustless setting. Wickström et al. (Wickström et al., 2022) propose the concept of AI market which builds a ML analysis network for IoT devices based on Ethereum. Similar to DMoE, Bittensor (Steeves et al., 2022) encapsulates neural network models as services and rewards peers contributing information-theoretic value to the system with TAO coins. Remote peers can be reused, for example, in the form of a MoE, or they can be pipelined to a new service.
The advantages of these blockchain-based systems are that transactions are transparent and workers can easily be incentivized. Therefore, it is a solution when it is challenging to establish a trust relationship between platform users. However, if available, we can replace these DAOs’ infrastructural components with a trusted crowdsourcing platform. Like Amazon Mechanical Turk, the incentivized mechanisms is no nessary for contract-based FL platforms as the rewards are set by employers in contract. While many studies consider Shapley value to evaluate the contribution of clients in FL systems (Zhan et al., 2021), we emphasize that a third-party evaluation can also be conducted in a crowdsourcing manner. Similarly, as the model community in query-based FL, contract platforms can further provide IP protection services to users (ref. Section 4.4). Dynamic fingerprints of models can be recorded and used for plagiarism detection. Three kinds of interests should be considered: 1) Interest of employers, to prevent the information leakage of task configurations; 2) Interest of workers, requiring fair evaluation of labor; 3) Interest of the public, to curb the infringement of privacy. However, this is not only a challenge from a technical perspective but also necessitates corresponding legislative oversight to provide guidelines to standardize this kind of AI marketplace commercialization. We believe that a comprehensive regulatory framework is essential to ensure data privacy, and establish fair practices in the evolving landscape of open FL platforms.
6. Conclusion
@@
Traditional federated learning systems with a server-dominated cooperation framework suppress the enthusiasm of participants and limit the further extension of such collaboration. To explore the opportunity to establish a more open and reciprocal cooperation platform, we investigate current progress in federated learning, decentralized machine learning, blockchain-based systems, and model reusing systems. In this way, we depict two rough sketches of open federated learning platforms: query-based federated learning and contract-based federated learning. Based on these two proposed platforms, we survey their possible supported techniques and their related legal issues, including ML licensing and copyrightability. We believe this survey can encourage a rethinking of current collaborative ML systems design and lead to the pervasive availability of AI for everyone.
Acknowledgements.

### 29 Feb 2024 v3

@@ VII-A Overiew
DeDES [190] Model Subset Model
@@
FedProto [155, 154] Class
@@ VII-B How to Design ML Microtasks
When we come to the scenario of contract-based FL, we first need to define how to design ML microtasks for crowd workers. Unfortunately, previous collaborative ML studies [25, 33] rarely discuss this question because they default to assuming that all employers or workers have the same resource type to fit into their modeling frameworks. On the contrary, in contract-based FL, we leave the freedom of design microtasks and choosing modeling method to the employers.
@@
The answer depends on what resources the employers have and what tasks the workers can perform. For example, in horizontal FL, the server should provide an initial model, and clients should have training data under the same feature space and provide computational power. However, in vertical FL, the clients should have training data under the same sample space (i.e., ID space). This difference determines which modeling framework employers can adopt. Following this idea, we provide comparisons of typical FL [8], distributed ML [224], and blockchain-based [221] modeling frameworks in TABLE VI. For simplicity, we introduce the Platform between Employer and Worker in some frameworks to make it adaptable for contract-based FL. You can quickly restore the original structure by merging Employer and Platform columns. To represent the transmission of resources, we mark resources provided by each group of entities with different colors. The computational resource used for training is marked with $\bigstar$. By this way, we can conveniently find the feasible and privacy-preserving modeling frameworks for collaborative learning. Here, we briefly introduce these frameworks group by scenarios [^5] From the viewpoint of Employers..
VII-B1 Train from scratch with workers’ private data
@@
There are four frameworks that support this scenario. The first is Centralized ML, which requires workers to upload their data to a cloud platform for modeling. However, this approach is not feasible for FL scenarios due to privacy concerns. The second is FedAvg [8], which requires workers to upload computed gradients upon the scratch model and their private data (horizontally split). Computing resources need to be provided by workers as well. The third is SplitNN [183], where each worker trains the cut layers of the model using their vertically split data. The hidden representations from these layers are uploaded to the platform for the training of the remaining layers. The last and distinctive is a Web3 system named Ocean [221]. It enables workers to upload only the metadata of their local dataset to a blockchain-based platform to attract buyers. Interested buyers then need to purchase datatokens to access the local dataset and deploy their modeling algorithm. To ensure the privacy and IP security of workers, only trusted algorithms are allowed, and only model predictions will be sent to buyers.
VII-B2 Train from FMs
@@
FMs, trained on larger-scale data with robust generalization abilities across various downstream tasks, serve as a solid foundation for collaborative training. Yuan et al. proposed DT-FM [214] to establish an effective geo-distributed learning system across 8 regions for training the language model GPT-3. Similarly, Borzunov et al.[223] recruited volunteer nodes to train a transformer model over the Internet. To enhance system robustness, Moshpit SGD [224] divides worker nodes into small, independent groups to ensure that all-reduce is not affected by the failure of a single participant. Additionally, VC-ASGD[225] divides the deep learning training job into asynchronous parameter update subtasks to improve scalability. In these cases, employers are required to supply the initial model and training data needed to launch training subtasks. This training paradigm is particularly suitable for worker nodes that possess computational and communication resources but lack their own training data.
@@
Instead of initiating training from scratch and consuming substantial computational resources, we can also collaboratively adapt these FMs to specific local tasks based on relatively restricted hardware and data size through fine-tuning. Recently, Woisetschläger et al. [229] have explored the opportunity to fine-tune large language models, such as FLAN-T5, by edge devices. Simultaneously, FS-LLM [226] employs Parameter-Efficient Fine-Tuning (PEFT) methods for federated fine-tuning of LLaMA-7B, minimizing communication and computation costs. FATE-LLM [230] offers a solution called FedHeteroLLM, leveraging KD to train a mentee model from its local pre-trained LLM for federated aggregation. To alleviate the computational burden associated with backpropagation-based optimization methods, FedKSeed [227] employs zeroth-order optimization (ZOO). Only seed and scalar gradients need to be transmitted for federated fine-tuning.
VII-B3 Ensemble workers’ private model
As we presented in §V, ensembling is an effective method to integrate knowledge, and the distributed version is proposed by DMoE [222]. It employs a Distributed Hash Table (DHT) to store metadata and worker statuses, constructing a decentralized expert network. To make inferences using this network, each model user must train a local gating network to select a subset of experts tailored to their input. No additional training required, Wang et al. purposed DeDES [190], which selects a diverse subsect of weak models from population and make inference by voting. The unique advantage of ensembling lies in its inherent support for model heterogeneity, and its integration is extensible. This capability enables the establishment of a cooperative network while maintaining flexibility and availability.
VII-B4 Aggregated workers’s private representations
@@
Another popular scheme for organizing ML microtasks involves sharing the learned representations of workers. For instance, the use of class-conditional average of last hidden layer activations (aka. Prototypes) can enhance class discrimination across different clients (e.g., FedProto [154, 155] and [228]). With no central server required, CCL [153] presents a decentralized learning approach based on cross-workers prototypes sharing. To configure the microtasks, employers only need to set the target class (no scratch model reqiured), and the platform can then collect and distribute per-class prototypes among workers.
VII-C Limitations of Contract-based FL
In contract-based FL systems, employers can publish their personalized ML crowdsourcing tasks, considering both their local resources and the resources of the target clients. However, to meet the flexibility requirements of widely supporting training methods, the contract platform should offer highly compatible APIs that can integrate various distributed training frameworks. Additionally, it needs to provide comprehensive audit against malicious users and plagiarism, significantly increasing the complexity compared to traditional FL and query-based FL. Meanwhile, to establish confidence among users, a fair and consensual third-party model evaluation mechanism should be established, which, however, is almost unexplored in the current academic studies.
VIII Conclusion
@@
Traditional federated learning systems with a server-dominated cooperation framework suppress the enthusiasm of participants and limit the further extension of such collaboration. To explore the opportunity to establish a more open and reciprocal cooperation platform, we investigate current progress in federated learning, decentralized machine learning, and model reusing systems. In this way, we depict two rough sketches of open federated learning platforms: query-based federated learning paltform and contract-based federated learning paltform. Based on these two proposed platforms, we survey their possible supported techniques and their related legal issues, including ML licensing and copyrightability. We believe this survey can encourage a rethinking of current collaborative ML systems design and lead to the pervasive availability of AI for everyone.
Acknowledgments
This research is supported by the National Research Foundation Singapore and DSO National Laboratories under the AI Singapore Programme (AISG Award No: AISG2-RP-2020-018). Any opinions, findings and conclusions or recommendations expressed in this material are those of the authors and do not reflect the views of National Research Foundation, Singapore.